OpenStudy (bossimbacon):

How many solutions does the equation have? 6-a=1/3(9a-18) A. one solution B. infinite solutions C. no solution

OpenStudy (bossimbacon):

@cj49 can you plz help me

OpenStudy (cj49):

A

OpenStudy (bossimbacon):

help me with one more plz @cj49

OpenStudy (cj49):

sure mate

OpenStudy (bossimbacon):

How many solutions does the equation have? 4x+1=5+2(2x-4) A. one solution B. infinite solutions C. no solution

OpenStudy (bossimbacon):

@cj49

OpenStudy (cj49):

C

OpenStudy (bossimbacon):

Which expressions would complete this equation so that it has one solution? 4(2x – 3) – 4x = Choose exactly two answers that are correct. A. 4x – 12 B. 3x – 4 C. 3(2x – 6) D. 4(x – 5)

OpenStudy (bossimbacon):

Which expressions would complete this equation so that it has one solution? 4(2x - 3) – 4x = Choose exactly two answers that are correct. A. 4x - 12 B. 3x – 4 C. 3(2x - 6) D. 4(x - 5)

OpenStudy (bossimbacon):

Which expressions would complete this equation so that it has one solution? 4(2x - 3) - 4x = Choose exactly two answers that are correct. A. 4x - 12 B. 3x - 4 C. 3(2x - 6) D. 4(x - 5)